OpenStudy (chrisplusian):

If I drop a 9v battery that is to pass two sensors spaced 4 inches apart, and I am holding the battery about two inches above the first sensor before I drop it what speed range could I expect in MPH for the battery to be going? I only need to know a range like 3-7MPH, or 10-15MPH, etc.

OpenStudy (agent0smith):

You could work it out, and convert it to mph. You can use physics formulas. \[\large v^2 = v_0^2 + 2a \Delta x\] initial speed would be 0, a=9.8 m/s^2, x = 4 inches but convert that to meters... about 0.1m

OpenStudy (agent0smith):

Sorry, 6 inches not 4

OpenStudy (agent0smith):

3-7MPH is reasonable

OpenStudy (chrisplusian):

I am not saying it would be 3-7 MPH, I am actually programming a microcontroller to measure speed of a falling object similar to a 9V battery in weight, size, and shape. SO I just wanted to verify my results. The cacluations I getting from the microcontroller says its travelling 11 MPH which seems wrong to me.

OpenStudy (agent0smith):

That is much too fast

OpenStudy (agent0smith):

Maximum speed for an ideal object (zero air resistance) after falling six inches is 3.87 mph

OpenStudy (chrisplusian):

Let me ask you this to verify my conversion factor.... if you went from inches per millisecond to MPH what would the conversion factor be? I got 227.2727. So when I multiplied my value which was given in inches per milliseconds by 227.2727 I believe it was then in MPH.

OpenStudy (agent0smith):

Oh god those are such terrible units. Divide the inches by 12 inches*5280 feet Multiply by 1000 milliseconds * 3600 seconds per hour

OpenStudy (agent0smith):

= 56.82

OpenStudy (agent0smith):

Googling inches per millisecond to miles per hour confirms my result. So there's your problem.

OpenStudy (chrisplusian):

\[\frac{ inch }{ millisecond }*\frac{ 1000millisecond }{ 1second }*\frac{ 3600second }{ 1 hour }*\frac{ 1foot }{12inches }*\frac{ 1mile }{ 5280ft }=\frac{ inch }{ millisecond }*227.2727\]

OpenStudy (chrisplusian):

Apparently I am off by a factor of four then.... don't know how that happened. Thanks for your help!

OpenStudy (agent0smith):

That thing looks right